An angular contact bearing of a lathe spindle is subjected to a cyclic radial load

during a manufacturing process in an automated plant. During the cycle, the

bearing experiences a radial load of 95 N for 30 s at 2 500 rpm, 75 N for 45 s at

3 500 rpm, and 115 N at 3 100 rpm for 15 s. Assume a shaft diameter of 15 mm.

Specify an appropriate bearing that will be used for 4 000 hours life with a 98 %

reliability. Give the bearing code as well as the relative ratings and dimensions.


1
Expert's answer
2021-04-28T07:25:21-0400



Equivalent load (Pe)


"_3\sqrt{\dfrac{N_1P_1^3 + N₂P₂³ +N_3 P_3 ³}{



N_1+N_2 + N_3}}"


"_3\sqrt{\dfrac{1250×95³ +2625 ×75³ + 775 ×115³}{



1250 + 2625 + 775}}"



Pe = 89.71N (Equivalent load/ mean cubic load)



Life of component = 4000h.


L10 Life = 6ON L10H/10⁶


N = N1 + N2 + N3 = 1250 + 2625 +775


N = 4650


+ ²N+IN = N


L10 Life = 60 × 4650 × 4000/10⁶


= 1116 milion revolutions


so, dynamic Load carrying capacity; (c)


c = Pe (L10)⅓


c = 89.71 (1116)⅓ = 930.52N